# Non-standard analysis for coherent risk estimation: hyperfinite representations, discrete Kusuoka formulae, and plug-in asymptotics

### v1

@@ 3.1. Standing conventions and the Loeb model
Remark 3.4 (Properties used in this paper).
@@ 3.7. The hyperfinite strong law and Glivenko–Cantelli theorem
Since $\|F_{N}-F\|_{\infty}\leqslant\varepsilon_{0}$ and $\varepsilon_{0}$ is infinitesimal, for any hyperreal $\alpha\in{}^{*}(0,1)$ with $\alpha>\varepsilon_{0}$ and $\alpha<1-\varepsilon_{0}$,
@@
The quantile shadow on $G=I_{N}\setminus D$. Take $k\notin D$, so $\alpha_{k}$ is not in the halo of any discontinuity of $q$ nor in the halo of $\{0,1\}$. This means $t:=\operatorname{st}(\alpha_{k})\in(0,1)\setminus\mathrm{Disc}(q)$, hence $q$ is continuous at $t$.
@@
hence $X_{k:N}\approx q(t)=q(\operatorname{st}(\alpha_{k}))\approx q(\alpha_{k})$ (using continuity again). Thus $k\in G$, and we conclude $G\supseteq I_{N}\setminus D$, hence $L(\mu_{N})(G)=1$.
@@
For the second term: $|q_{N}-q_{N}^{(M)}|=|q_{N}|\mathbf{1}_{|q_{N}|>M}$. Since $\mathsf{E}|X_{1}|<\infty$, the hyperfinite SLLN gives $\frac{1}{N}\sum_{k=1}^{N}|X_{k}|\approx\mathsf{E}|X_{1}|<\infty$. By the standard integration identity for order statistics, $\int_{0}^{1}|q_{N}|\,d\alpha=\frac{1}{N}\sum_{k=1}^{N}|X_{k:N}|$, which has finite standard part. Thus $\int_{0}^{1}|q_{N}-q_{N}^{(M)}|\to 0$ as $M\to\infty$ (in the sense that for any standard $\delta>0$, there exists standard $M$ such that this integral $\delta$).
For the third term: $\int_{0}^{1}|q-q^{(M)}|\to 0$ as $M\to\infty$ by $\mathsf{E}|X_{1}|<\infty$.
Given $\delta>0$, choose $M$ large enough that the second and third terms are each $\delta/3$. Then choose the infinitesimal such that the first term is infinitesimal. Taking standard parts, $\operatorname{st}(\int_{0}^{1}|q_{N}-q|)\leqslant 2\delta/3<\delta$. Since $\delta$ was arbitrary, $\int_{0}^{1}|q_{N}-q|\approx 0$.
@@ 4.1. The hyperfinite dictionary
Moreover, if $X\in L^{\infty}$ has a bounded internal lifting $\tilde{X}$ with $|\tilde{X}|\leqslant M$, then for any $Z\in{}^{*}\mathcal{Z}$ the product $\tilde{X}\cdot Z$ is $S$-integrable and
$\mathsf{E}[-XZ]=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big).$
@@
For the product: since $|\tilde{X}|\leqslant M$ and $Z\geqslant 0$ is $S$-integrable, $|\tilde{X}\cdot Z|\leqslant M\cdot Z$ is dominated by an $S$-integrable function, hence $\tilde{X}\cdot Z$ is $S$-integrable. ∎
@@ 4.2. Hyperfinite robust representation
where $^{*}\mathcal{Z}$ is the internal non-standard extension of $\mathcal{Z}$, and elements $Z\in{}^{*}\mathcal{Z}$ are $S$-integrable by Lemma 4.2. The supremum over $^{*}\mathcal{Z}$ is internal, and is attained by some $Z^{\sharp}\in{}^{*}\mathcal{Z}$.
@@
The attainment statement follows from transfer: the standard maximiser $Z^{*}\in\mathcal{Z}$ extends to $Z^{\sharp}={}^{*}Z^{*}\in{}^{*}\mathcal{Z}$, and $a^{\sharp}:=\Psi(Z^{\sharp})$ achieves a value infinitely close to the supremum over $\mathcal{A}_{N}$. ∎
@@ 4.3. From hyperfinite representation to CRE representation
Proof of Theorem 2.11 (finite Choquet / distortion representation).
Let $\hat{\rho}_{n}$ be coherent, law-invariant, and comonotonic additive. By Theorem 2.10 there exists a non-empty convex set $M^{s}_{\hat{\rho}_{n}}\subseteq\Delta_{n}^{\downarrow}$ such that
$\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i:n}).$
A standard representation theorem for comonotonic additive, monotone functionals on a finite lattice states that such a functional is a Choquet integral with respect to a unique capacity (called the core of the functional); see, e.g., [9, Ch. 5]. In the present finite-dimensional setting, one may take the capacity $c:2^{\{1,\dots,n\}}\to[0,1]$ defined by
$c(A):=\hat{\rho}_{n}(-\mathbf{1}_{A}),\qquad A\subseteq\{1,\dots,n\},$
where $\mathbf{1}_{A}\in\mathbb{R}^{n}$ is the indicator vector of $A$.
Law invariance implies that $c(A)$ depends only on $|A|$, hence $c$ is determined by a unique distortion function on the grid $\{0,1,\dots,n\}$:
$g(k):=c(\{1,2,\dots,k\})=\hat{\rho}_{n}(-\mathbf{1}_{\{1,\dots,k\}}),\qquad k=0,\dots,n,$
with $g(0)=0$ and $g(n)=1$. Coherence (in particular subadditivity) is equivalent in this finite setting to discrete concavity of $g$ (equivalently, to $\Delta_{n}^{\downarrow}$-monotonicity of the resulting weights).
By the finite Choquet integral formula (see, e.g., [9, Ch. 5] or the risk-measure presentation in [11]), for every $x\in\mathbb{R}^{n}$,
$\hat{\rho}_{n}(x)=\sum_{i=1}^{n}\bigl(g(i)-g(i-1)\bigr)\,(-x_{i:n}).$
Setting $a_{i}:=g(i)-g(i-1)$ yields $a\in\Delta_{n}^{\downarrow}$. Uniqueness follows because the values $\hat{\rho}_{n}(-\mathbf{1}_{\{1,\dots,k\}})=g(k)$ recover the increments $a_{i}=g(i)-g(i-1)$, hence the weight vector is uniquely determined.
Consequently $M^{s}_{\hat{\rho}_{n}}$ must be the singleton $\{a\}$, proving the theorem. ∎
@@ 7.2. Spectral L-estimator consistency theorem
almost surely, and the right-hand side vanishes as $M\to\infty$. Hence $\int_{0}^{1}|q_{n}-q_{X}|\,d\alpha\to 0$ almost surely, and therefore
@@ 8.2. Consistency theorem
Assumption (K1) excludes risk measures that put mass arbitrarily close to $\alpha=0$, which is necessary because $\operatorname{ES}_{\alpha}(X)\to$-$\infty$ as $\alpha\downarrow 0$ for unbounded-below $X$, and discretisation near 0 is unstable. A typical choice for $\widehat{\operatorname{ES}}_{\alpha,n}$ is the discrete ES $\mathrm{dES}_{\lceil n\alpha\rceil/n}$ from (13), for which (K2)–(K3) hold; see Theorem 7.2.
@@
Fix a bounded random variable $X$ and an i.i.d. sample $(X_{i})_{i\geqslant 1}$ with law $X$. For $\nu\in\mathcal{M}$ set
$I(\nu):=\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha),\qquad\hat{I}_{n}(\nu):=\sum_{i=1}^{n}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}]).$
Then $\rho(X)=\sup_{\nu\in\mathcal{M}}I(\nu)$ and $\hat{\rho}_{n}(X_{1},\ldots,X_{n})=\sup_{\nu\in\mathcal{M}}\hat{I}_{n}(\nu)$. Since $|\sup_{\nu}a(\nu)-\sup_{\nu}b(\nu)|\leqslant\sup_{\nu}|a(\nu)-b(\nu)|$, it suffices to prove that
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\longrightarrow 0\qquad\text{almost surely.}$
Fix $\varepsilon>0$ and choose $\delta\in(0,1)$ such that $\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\varepsilon$ (assumption (K1)). For each $\nu\in\mathcal{M}$, split
$I(\nu)=\int_{(0,\delta]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)+\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)=:I_{0,\delta}(\nu)+I_{\delta,1}(\nu),$
and split $\hat{I}_{n}(\nu)$ analogously into the terms with $\alpha_{i,n}\leqslant\delta$ and $\alpha_{i,n}>\delta$:
$\hat{I}_{n}(\nu)=\sum_{\alpha_{i,n}\leqslant\delta}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}])\\ +\sum_{\alpha_{i,n}>\delta}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}])=:\hat{I}_{0,\delta,n}(\nu)+\hat{I}_{\delta,1,n}(\nu).$
Because $X$ is bounded we have $|\operatorname{ES}_{\alpha}(X)|\leqslant\|X\|_{\infty}$ for all $\alpha\in(0,1]$. Therefore,
$\sup_{\nu\in\mathcal{M}}|I_{0,\delta}(\nu)|\leqslant\|X\|_{\infty}\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\|X\|_{\infty}\,\varepsilon.$
Moreover, since $X$ is bounded, $\|(X_{1},\ldots,X_{n})\|_{\infty}\leqslant\|X\|_{\infty}$ almost surely, and assumption (K3) yields $|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})|\leqslant C\|X\|_{\infty}$ for all $\alpha,n$ on an almost sure event. Hence,
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{0,\delta,n}(\nu)|\leqslant C\|X\|_{\infty}\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant C\|X\|_{\infty}\,\varepsilon\qquad\text{almost surely.}$
Combining these two bounds gives
$\sup_{\nu\in\mathcal{M}}\bigl|\hat{I}_{0,\delta,n}(\nu)-I_{0,\delta}(\nu)\bigr|\leqslant(1+C)\|X\|_{\infty}\,\varepsilon\qquad\text{almost surely.}$
It remains to control the contribution from $(\delta,1]$. Write
$\hat{I}_{\delta,1,n}(\nu)-I_{\delta,1}(\nu)$ $=\sum_{\alpha_{i,n}>\delta}\Big(\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha_{i,n}}(X)\Big)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])$
$\quad+\sum_{\alpha_{i,n}>\delta}\operatorname{ES}_{\alpha_{i,n}}(X)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])-\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha).$
For the first term we simply use $\nu((\delta,1])\leqslant 1$ and assumption (K2):
@@
$\sup_{\nu\in\mathcal{M}}\left|\sum_{\alpha_{i,n}>\delta}\Big(\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha_{i,n}}(X)\Big)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])\right|\\ \leqslant\sup_{\alpha\in[\delta,1]}\left|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha}(X)\right|\longrightarrow 0$
almost surely.
For the second term, define $g(\alpha):=\operatorname{ES}_{\alpha}(X)$ on $\delta,1]$. Since $X$ is bounded, $q_{X}$ is bounded, so $\alpha\mapsto\int_{0}^{\alpha}q_{X}(u)\,du$ is continuous and therefore $g(\alpha)=-(1/\alpha)\int_{0}^{\alpha}q_{X}(u)\,du$ is continuous on $\delta,1]$. By compactness, $g$ is uniformly continuous on $\delta,1]$. Let $\omega_{g}(h)$ denote its modulus of continuity:
$\omega_{g}(h):=\sup\{|g(\alpha)-g(\beta)|:\ \alpha,\beta\in[\delta,1],\ |\alpha-\beta|\leqslant h\}.$
Let $\Delta_{n}:=\max_{i}(\alpha_{i,n}-\alpha_{i-1,n})$, which tends to 0 by (K4). Then, for every $\nu\in\mathcal{M}$,
$\left|\sum_{\alpha_{i,n}>\delta}\operatorname{ES}_{\alpha_{i,n}}(X)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])-\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)\right|$
$\qquad=\left|\sum_{\alpha_{i,n}>\delta}\int_{(\alpha_{i-1,n},\alpha_{i,n}]}\bigl(g(\alpha_{i,n})-g(\alpha)\bigr)\,\nu(d\alpha)\right|$
$\qquad\leqslant\sum_{\alpha_{i,n}>\delta}\int_{(\alpha_{i-1,n},\alpha_{i,n}]}\omega_{g}(\Delta_{n})\,\nu(d\alpha)\leqslant\omega_{g}(\Delta_{n}).$
This bound is uniform in $\nu$, and $\omega_{g}(\Delta_{n})\to 0$ as $n\to\infty$.
Putting everything together, we obtain on an almost sure event:
$\limsup_{n\to\infty}\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\leqslant(1+C)\|X\|_{\infty}\,\varepsilon.$
Since $\varepsilon>0$ is arbitrary, $\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\to 0$ almost surely, we conclude convergence $\hat{\rho}_{n}(X_{1},\ldots,X_{n})\to\rho(X)$ a.s. ∎

### v2

@@ 3.1. Standing conventions and the Loeb model
Remark 3.4.
@@ 3.7. The hyperfinite strong law and Glivenko–Cantelli theorem
Since $\|F_{N}-F\|_{\infty}\leqslant\varepsilon_{0}$ and $\varepsilon_{0}$ is infinitesimal, for any hyperreal $\alpha\in{}^{*}(0,1)$ with $\alpha>\varepsilon_{0}$ and $\alpha<1-\varepsilon_{0}$, writing $q_{N}(\alpha):=X_{\lceil N\alpha\rceil:N}$ for the internal empirical quantile function,
@@
The quantile shadow on $G=I_{N}\setminus D$. Take $k\notin D$ with $k<N$, so $\alpha_{k}$ is not in the halo of any discontinuity of $q$ nor in the halo of $\{0,1\}$. This means $t:=\operatorname{st}(\alpha_{k})\in(0,1)\setminus\mathrm{Disc}(q)$, hence $q$ is continuous at $t$.
@@
hence $X_{k:N}\approx q(t)=q(\operatorname{st}(\alpha_{k}))\approx q(\alpha_{k})$ (using continuity again). Thus $k\in G$, and we conclude $G\supseteq(I_{N}\setminus D)\cap\{k:k<N\}$, hence $L(\mu_{N})(G)=1$ (since $\{N\}$ has Loeb measure $1/N\approx 0$).
@@
For the second term: note that
$\int_{0}^{1}|q_{N}-q_{N}^{(M)}|\,d\alpha=\frac{1}{N}\sum_{i=1}^{N}|X_{i}|\,\mathbf{1}_{\{|X_{i}|>M\}},$
because $q_{N}$ is the empirical quantile step function. Apply Theorem 3.18 to the integrable variables
$Y_{i}^{(M)}:=|X_{i}|\,\mathbf{1}_{\{|X_{i}|>M\}}.$
Then
$\int_{0}^{1}|q_{N}-q_{N}^{(M)}|\,d\alpha\approx\mathsf{E}\!\left[|X_{1}|\,\mathbf{1}_{\{|X_{1}|>M\}}\right].$
Since $X_{1}\in L^{1}$, the right-hand side tends to 0 as $M\to\infty$.
For the third term,
$\int_{0}^{1}|q-q^{(M)}|\,d\alpha=\mathsf{E}\!\left[|X_{1}|\,\mathbf{1}_{\{|X_{1}|>M\}}\right]\xrightarrow[M\to\infty]{}0.$
Hence, for every standard $\delta>0$, one can choose standard $M$ so large that the second and third terms are both $\delta/3$, while for this fixed $M$ the first term is infinitesimal. Therefore
$\int_{0}^{1}|q_{N}-q|\,d\alpha\approx 0.$
@@ 4.1. The hyperfinite dictionary
Moreover, if $X\in L^{\infty}$ has a bounded internal lifting $\tilde{X}$ with $|\tilde{X}|\leqslant M$, and if $f_{X}:\mathcal{Z}\to\mathbb{R}$ is defined by
$f_{X}(Z_{0}):=\mathsf{E}[-XZ_{0}],\qquad Z_{0}\in\mathcal{Z},$
then for any $Z\in{}^{*}\mathcal{Z}$ the product $\tilde{X}\cdot Z$ is $S$-integrable and
$^{*}f_{X}(Z)=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big).$
In particular, for standard $Z_{0}\in\mathcal{Z}$,
$f_{X}(Z_{0})=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,{}^{*}Z_{0}(k)\Big).$
@@
For the product: since $|\tilde{X}|\leqslant M$ and $Z\geqslant 0$ is $S$-integrable, we have
$|\,\tilde{X}\cdot Z\,|\leqslant MZ,$
so $\tilde{X}\cdot Z$ is $S$-integrable.
If $f_{X}(Z_{0}):=\mathsf{E}[-XZ_{0}]$ on $\mathcal{Z}$, then for $Z\in{}^{*}\mathcal{Z}$ the non-standard extension $^{*}f_{X}(Z)$ is represented on the Loeb model by the $S$-integrable internal product $(-\tilde{X})\cdot Z$. Therefore
$^{*}f_{X}(Z)=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big),$
as claimed. ∎
@@ 4.2. Hyperfinite robust representation
where $^{*}\mathcal{Z}$ is the internal non-standard extension of $\mathcal{Z}$, and elements $Z\in{}^{*}\mathcal{Z}$ are $S$-integrable by Lemma 4.2. Writing $f_{X}(Z_{0}):=\mathsf{E}[-XZ_{0}]$ for $Z_{0}\in\mathcal{Z}$, the internal dual problem
$\sup_{Z\in{}^{*}\mathcal{Z}}{}^{*}f_{X}(Z)$
is attained by some $Z^{\sharp}\in{}^{*}\mathcal{Z}$ (for instance $Z^{\sharp}={}^{*}Z^{*}$, where $Z^{*}\in\mathcal{Z}$ is any standard maximiser of $f_{X}$). For the hyperfinite-sum functional
$S_{X}(Z):=\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k),$
the same $Z^{\sharp}$ is a near-maximiser in the sense that
$\operatorname{st}\bigl(S_{X}(Z^{\sharp})\bigr)=\rho(X)=\operatorname{st}\Big(\sup_{Z\in{}^{*}\mathcal{Z}}S_{X}(Z)\Big).$
@@
Thus $Z^{\sharp}={}^{*}Z^{*}$ attains the exact maximum of the internal functional $^{*}f_{X}$, and $a^{\sharp}:=\Psi(Z^{\sharp})$ yields a value whose standard part equals the supremum over $\mathcal{A}_{N}$. In general we only claim near-attainment for the hyperfinite-sum functionals. ∎
@@ 4.3. From hyperfinite representation to CRE representation
Proof of Theorem 2.11.
The implication $\Leftarrow$” is immediate from Theorem 2.10: if
$\hat{\rho}_{n}(x)=\sum_{i=1}^{n}a_{i}(-x_{i:n}),\qquad a\in\Delta_{n}^{\downarrow},$
then $\hat{\rho}_{n}$ is a law-invariant CRE. To see comonotonicity, let $x,y\in\mathbb{R}^{n}$ be comonotonic. Then there exists a permutation $\pi$ such that both $\pi(x)$ and $\pi(y)$ are increasing. Consequently,
$s(x+y)=s(x)+s(y),$
and therefore
$\hat{\rho}_{n}(x+y)=\sum_{i=1}^{n}a_{i}\bigl(-(x+y)_{i:n}\bigr)=\sum_{i=1}^{n}a_{i}(-x_{i:n})+\sum_{i=1}^{n}a_{i}(-y_{i:n})=\hat{\rho}_{n}(x)+\hat{\rho}_{n}(y).$
For the converse, let $M^{s}_{\hat{\rho}_{n}}\subseteq\Delta_{n}^{\downarrow}$ be a representing set from Theorem 2.10, and define
$K_{\hat{\rho}_{n}}:=\overline{M^{s}_{\hat{\rho}_{n}}}^{\,\mathbb{R}^{n}}\subseteq\Delta_{n}^{\downarrow}.$
Since the map $a\mapsto\langle a,-s(x)\rangle$ is continuous for each fixed $x$, passing from $M^{s}_{\hat{\rho}_{n}}$ to $K_{\hat{\rho}_{n}}$ does not change the support function:
$\hat{\rho}_{n}(x)=\sup_{a\in K_{\hat{\rho}_{n}}}\langle a,-s(x)\rangle,\qquad x\in\mathbb{R}^{n}.$
Moreover, $K_{\hat{\rho}_{n}}$ is compact. For $x\in\mathbb{R}^{n}$, define
$N(x):=\{a\in K_{\hat{\rho}_{n}}:\ \hat{\rho}_{n}(x)=\langle a,-s(x)\rangle\}.$
Then each $N(x)$ is non-empty and closed in the compact set $K_{\hat{\rho}_{n}}$.
We claim that the family $\{N(x):x\in\mathbb{R}^{n}\}$ has the finite intersection property. Fix $x^{1},\dots,x^{K}\in\mathbb{R}^{n}$ and set
$y:=\sum_{j=1}^{K}s(x^{j}).$
Since each $s(x^{j})$ is increasing, the vectors $s(x^{1}),\dots,s(x^{K})$ are pairwise comonotonic. Hence, by comonotonicity and law invariance,
$\hat{\rho}_{n}(y)=\sum_{j=1}^{K}\hat{\rho}_{n}(s(x^{j}))=\sum_{j=1}^{K}\hat{\rho}_{n}(x^{j}).$
Choose $a\in N(y)$. Then
$\hat{\rho}_{n}(y)=\langle a,-y\rangle=\sum_{j=1}^{K}\langle a,-s(x^{j})\rangle.$
Because $a\in K_{\hat{\rho}_{n}}$, we have
$\langle a,-s(x^{j})\rangle\leqslant\hat{\rho}_{n}(x^{j}),\qquad j=1,\dots,K,$
since $\hat{\rho}_{n}(x^{j})=\sup_{b\in K_{\hat{\rho}_{n}}}\langle b,-s(x^{j})\rangle$. Comparing the last two displays forces equality term by term:
$\langle a,-s(x^{j})\rangle=\hat{\rho}_{n}(x^{j}),\qquad j=1,\dots,K.$
Thus
$a\in\bigcap_{j=1}^{K}N(x^{j}),$
which proves the finite intersection property.
Since $K_{\hat{\rho}_{n}}$ is compact,
$\bigcap_{x\in\mathbb{R}^{n}}N(x)\neq\varnothing.$
Pick $a$ in this intersection. Then
$\hat{\rho}_{n}(x)=\langle a,-s(x)\rangle=\sum_{i=1}^{n}a_{i}(-x_{i:n}),\qquad x\in\mathbb{R}^{n}.$
Finally, to prove uniqueness, suppose $a,b\in\Delta_{n}^{\downarrow}$ both satisfy the representation. For
$x^{(k)}:=(-1,\dots,-1,0,\dots,0)\in\mathbb{R}^{n}$
with exactly $k$ entries equal to 1-1, we obtain
$\sum_{i=1}^{k}a_{i}=\hat{\rho}_{n}(x^{(k)})=\sum_{i=1}^{k}b_{i},\qquad k=1,\dots,n.$
Subtracting the relation for k1k-1 from that for $k$ gives $a_{k}=b_{k}$ for every $k$, so $a=b$. ∎
@@ 7.2. Spectral L-estimator consistency theorem
almost surely by the strong law, and the right-hand side vanishes as $M\to\infty$. Hence $\int_{0}^{1}|q_{n}-q_{X}|\,d\alpha\to 0$ almost surely, and therefore
@@ 8.2. Consistency theorem
Assumption (K1) excludes risk measures that put mass arbitrarily close to $\alpha=0$, which is necessary because $\operatorname{ES}_{\alpha}(X)\to$-$\infty$ as $\alpha\downarrow 0$ for unbounded-below $X$, and discretisation near 0 is unstable. A typical choice for $\widehat{\operatorname{ES}}_{\alpha,n}$ is the discrete ES $\mathrm{dES}_{\lceil n\alpha\rceil/n}$ from (13). In this case, (K3) is immediate from the definition, while (K2) is a uniform expected-shortfall consistency statement on $\delta,1]$ and requires a separate argument or a dedicated reference; see, for example, [7].
@@
Fix a bounded random variable $X$ and an i.i.d. sample $(X_{i})_{i\geqslant 1}$ with law $X$. For $I_{i,n}:=(\alpha_{i-1,n},\alpha_{i,n}]$ and $\nu\in\mathcal{M}$, write
$I(\nu):=\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha),\qquad\hat{I}_{n}(\nu):=\sum_{i=1}^{n}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu(I_{i,n}).$
Then
$\rho(X)=\sup_{\nu\in\mathcal{M}}I(\nu),\qquad\hat{\rho}_{n}(X_{1},\ldots,X_{n})=\sup_{\nu\in\mathcal{M}}\hat{I}_{n}(\nu).$
Since $|\sup_{\nu}a(\nu)-\sup_{\nu}b(\nu)|\leqslant\sup_{\nu}|a(\nu)-b(\nu)|$, it suffices to show
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\xrightarrow{\mathrm{a.s.}}0.$
Decompose
$\hat{I}_{n}(\nu)-I(\nu)=A_{n}(\nu)+B_{n}(\nu),$
where
$A_{n}(\nu):=\sum_{i=1}^{n}\Big(\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha_{i,n}}(X)\Big)\,\nu(I_{i,n}),$
and
$B_{n}(\nu):=\sum_{i=1}^{n}\operatorname{ES}_{\alpha_{i,n}}(X)\,\nu(I_{i,n})-\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha).$
Fix $\varepsilon>0$. By (K1), choose $\delta\in(0,1)$ such that
$\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\varepsilon.$
For the estimator-error term $A_{n}(\nu)$, using $|\operatorname{ES}_{\alpha}(X)|\leqslant\|X\|_{\infty}$ and (K3), we get
@@
$\sup_{\nu\in\mathcal{M}}|A_{n}(\nu)|$ $\leqslant\sup_{\alpha\in[\delta,1]}\left|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha}(X)\right|$
$\qquad+(C+1)\|X\|_{\infty}\sup_{\nu\in\mathcal{M}}\nu((0,\delta]).$
Hence, by (K2),
$\limsup_{n\to\infty}\sup_{\nu\in\mathcal{M}}|A_{n}(\nu)|\leqslant(C+1)\|X\|_{\infty}\,\varepsilon\qquad\text{a.s.}$
It remains to control $B_{n}(\nu)$. Define
$g(\alpha):=\operatorname{ES}_{\alpha}(X),\qquad\alpha\in(0,1].$
Because $X$ is bounded, the quantile function $q_{X}$ is bounded and monotone on $(0,1)$. Hence $q_{X}(0+):=\lim_{\alpha\downarrow 0}q_{X}(\alpha)$ exists and
$\lim_{\alpha\downarrow 0}\operatorname{ES}_{\alpha}(X)=-\lim_{\alpha\downarrow 0}\frac{1}{\alpha}\int_{0}^{\alpha}q_{X}(u)\,du=-q_{X}(0+).$
Thus $g$ extends continuously to $[0,1]$, and therefore is uniformly continuous there. Let $\omega_{g}$ be its modulus of continuity and let
$\Delta_{n}:=\max_{1\leqslant i\leqslant n}(\alpha_{i,n}-\alpha_{i-1,n}).$
Then, for every $\nu\in\mathcal{M}$,
$|B_{n}(\nu)|$ $=\left|\sum_{i=1}^{n}\int_{I_{i,n}}\bigl(g(\alpha_{i,n})-g(\alpha)\bigr)\,\nu(d\alpha)\right|$
$\leqslant\sum_{i=1}^{n}\int_{I_{i,n}}\omega_{g}(\Delta_{n})\,\nu(d\alpha)\leqslant\omega_{g}(\Delta_{n}).$
By (K4), $\Delta_{n}\to 0$, so $\omega_{g}(\Delta_{n})\to 0$. Hence
$\sup_{\nu\in\mathcal{M}}|B_{n}(\nu)|\xrightarrow[n\to\infty]{}0.$
Combining the two bounds, we obtain
$\limsup_{n\to\infty}\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\leqslant(C+1)\|X\|_{\infty}\,\varepsilon\qquad\text{a.s.}$
Since $\varepsilon>0$ is arbitrary, this proves
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\xrightarrow{\mathrm{a.s.}}0,$
and therefore $\hat{\rho}_{n}(X_{1},\ldots,X_{n})\to\rho(X)$ almost surely. ∎